Decide how the objects in the following table would be listed, if they were listed in order of increasing kinetic energy. That is, select 1 next to the object with the lowest kinetic energy, select 2 next to the object with the second lowest kinetic energy, and so on. object kinetic energy A 87. kg runner moving at 8.5 m/s.(choose one) A 104. kg runner moving at 9.3 m/s A 96. kg runner moving at 9.3 m/s. (choose one) A 96. kg runner moving at 8.5 m/s. (choose one) alo (ehoose one) Ar
Answer #1

for 1st row:
KE = 0.5*m*v^2
= 0.5*87*8.5^2
= 3143 J

for 2nd row:
KE = 0.5*m*v^2
= 0.5*104*9.3^2
= 4497 J

for 3rd row:
KE = 0.5*m*v^2
= 0.5*96*9.3^2
= 4152 J

for 4th row:
KE = 0.5*m*v^2
= 0.5*96*8.5^2
= 3468 J

Answer:
1
4
3
2